# Self-Critique and Refinement for Faithful Natural Language Explanations

### v1

@@ Abstract
With the rapid development of large language models (LLMs), natural language explanations (NLEs) have become increasingly important for understanding model predictions. However, these explanations often fail to faithfully represent the model’s actual reasoning process. While existing work has demonstrated that LLMs can self-critique and refine their initial outputs for various tasks, this capability remains unexplored for improving explanation faithfulness. To address this gap, we introduce Self-critique and Refinement for Natural Language Explanations (SR-NLE), a framework that enables models to improve the faithfulness of their own explanations – specifically, post-hoc NLEs – through an iterative critique and refinement process without external supervision. Our framework leverages different feedback mechanisms to guide the refinement process, including natural language self-feedback and, notably, a novel feedback approach based on feature attribution that highlights important input words. Our experiments across three datasets and four state-of-the-art LLMs demonstrate that SR-NLE significantly reduces unfaithfulness rates, with our best method achieving an average unfaithfulness rate of 36.02%, compared to 54.81% for baseline – an absolute reduction of 18.79%. These findings reveal that the investigated LLMs can indeed refine their explanations to better reflect their actual reasoning process, requiring only appropriate guidance through feedback without additional training or fine-tuning.
@@ 1 Introduction
With the rapid development of large language models (LLMs), both closed-source models (OpenAI et al., 2024; Gemini et al., 2025) and open-source alternatives (Qwen et al., 2025; Grattafiori et al., 2024) have demonstrated remarkable capabilities across a wide range of natural language processing (NLP) tasks. However, despite these advancements, understanding the reasoning behind their predictions remains a critical challenge – especially in applications that demand trust and accountability.
Natural Language Explanations (NLEs) have emerged as a promising solution by offering human-readable justifications for model predictions without requiring access to internal model mechanisms. Yet, ensuring their faithfulness remains a significant challenge. Recent studies have shown that NLEs generated by LLMs often fail to reflect the actual reasoning process of the model (Atanasova et al., 2023; Turpin et al., 2023; Lanham et al., 2023).
While prior work has primarily relied on changes of the model architecture or additional fine-tuning Yuan et al. (2025); Wang et al. (2023a), we instead explore whether models possess the capability to independently assess and refine their own explanations. Supporting this direction, recent studies have shown that LLMs are capable of improving their outputs through iterative self-refinement (Madaan et al., 2023; Shinn et al., 2023). Following this, we ask whether LLMs know if and when their NLEs are faithful to their own internal reasoning by providing self-critique and refining their NLEs for improved faithfulness.
Building on this idea, we propose Self-critique and Refinement for NLEs (SR-NLE), a framework that enables models to improve the faithfulness of their own explanations through an iterative critique and refinement process without external supervision, as only the model itself has access to its internal reasoning and is therefore best positioned to assess explanation faithfulness. Our framework specifically targets post-hoc NLEs, where the explanation is generated after the model makes a prediction. Starting from an initial explanation, the model receives feedback identifying potential issues and generates a refined explanation accordingly. This process can be repeated multiple times, enabling incremental improvements.
@@
  • We introduce SR-NLE, a novel framework that seamlessly enables models to improve the faithfulness of their explanations through iterative self-critique and refinement guided by different feedback mechanisms, without external assistance, architectural modifications or specialized training.
  • We propose and evaluate multiple faithfulness feedback strategies, including natural language feedback (NLF) and a novel feedback mechanism – Important Word Feedback (IWF), that leverages feature attribution to identify overlooked important input words in generated NLEs.
@@ 2 Related Works
Natural Language Explanations (NLEs) provide human-readable justifications for model predictions, traditionally obtained via supervised training on annotated datasets (Camburu et al., 2018; Rajani et al., 2019; Atanasova et al., 2020b). More recently, LLMs have enabled NLE generation via in-context learning (Brown et al., 2020). A prominent example is chain-of-thought reasoning (Wei et al., 2022; Kojima et al., 2022), where LLMs generate intermediate reasoning steps alongside the model prediction. In contrast, our work focuses on post-hoc NLEs, where explanations are generated after the model prediction is made. Despite these advances in NLE generation, numerous studies have identified a gap between generated NLEs and the model’s actual reasoning process (Atanasova et al., 2023; Turpin et al., 2023; Lanham et al., 2023). To quantify this faithfulness gap, researchers have proposed various automatic evaluation metrics, such as counterfactual tests (Atanasova et al., 2023; Siegel et al., 2024) and association-based measures (Wiegreffe et al., 2022; Parcalabescu and Frank, 2024). In this work, we adopt counterfactual tests (Atanasova et al., 2023) as our primary evaluation method, as they offer instance-level, automatic assessments of explanation faithfulness without requiring human annotations.
@@
Recent work has shown that LLMs can improve their own outputs via iterative self-refinement. The Self-Refine (Madaan et al., 2023) demonstrates that models can critique and revise their own outputs, leading to improved performance across a variety of tasks. Similarly, Shinn et al. (2023) explore self-reflection mechanisms for agent-level reasoning. While these works establish the general potential of self-improvement, they do not specifically address the challenge of improving explanation faithfulness. Building on the general idea of self-refinement, Cross-Refine (Wang et al., 2024) applies this paradigm to NLE generation. Their framework adopts a cross-model design, where one LLM generates the initial explanation and another, separate LLM provides feedback and suggestions for revision. In contrast, SR-NLE operates entirely within a single model, leveraging its internal capabilities for both critique and refinement. Furthermore, SR-NLE focuses specifically on improving faithfulness using automated counterfactual tests for objective evaluation. In comparison, Cross-Refine primarily evaluates explanation quality through multiple automated metrics, while relying on human judgments for faithfulness assessment.
@@
Input feature attribution methods quantify how much each input token contributes to a model’s prediction. Common approaches include Shapley values (Lundberg and Lee, 2017), attention weights (Jain and Wallace, 2019), and integrated gradients (Sundararajan et al., 2017). Traditionally, such methods are used as post-hoc explanations to help humans understand which parts of the input are most influential in model decision-making across various tasks. In this work, we explore a novel use of attribution in the context of NLE generation with LLMs. Rather than using attribution to explain model predictions to humans, we treat it as a feedback signal to guide explanation refinement. Specifically, we identify the most important input words based on attribution scores and use them to guide the model in generating more faithful NLEs.
@@ 3 Method
Figure 2: The SR-NLE framework. (a) Answer and Explanation Generation Phase: The framework produces the answer and initial explanation. (b) Iterative Critique and Refinement Phase: The framework iteratively improves explanations through feedback-refinement loops over multiple rounds.
In this section, we present SR-NLE, a framework for improving the faithfulness of NLEs generated by LLMs. SR-NLE employs an iterative self-critique and refinement process, enabling LLMs to progressively identify faithfulness issues in their own NLEs and make targeted improvements thereof. This framework leverages the in-context learning and self-improvement capabilities of large language models, without requiring human involvement or additional models for feedback.
@@ 3.1 Preliminary
The SR-NLE framework operates on the assumption that LLMs have the capability to identify and improve their own explanations when guided with appropriate prompts. Our framework relies entirely on a single model $\mathcal{M}$ for all components, without requiring human involvement or additional models. For an input $x$, the model first predicts an answer $y$ and produces an initial explanation $e^{0}$, then through an iterative process of self-critique and refinement, after each round $r$, we obtain a progressively improved explanation $e^{r}$. To direct the model in different stages of the framework, we employ four categories of prompts: $p_{\text{ans}}$ for answer generation, $p_{\text{exp}}$ for explanation generation, $p_{\text{fb}}$ for feedback generation, and $p_{\text{ref}}$ for refinement generation, where both feedback and refinement prompts have two variants corresponding to our two feedback approaches: natural language feedback (NLF) and important word feedback (IWF). Throughout the framework, we use $\oplus$” to denote filling a prompt template with its variables.
@@ 3.2 SR-NLE Framework
Our SR-NLE framework consists of two main phases: (a) Answer and Explanation Generation, which produces the answer and its initial explanation, and (b) Iterative Critique and Refinement, which progressively improves the explanation through multiple rounds. Figure 2 illustrates this two-phase process. The algorithmic formulation of the framework is provided as Algorithm 1.
@@ 3.2.2 Iterative Critique and Refinement
For each refinement round $r$, the model generates feedback on the current explanation. We explore two distinct feedback approaches:
  • Natural-Language Feedback (NLF). With this approach, $\mathcal{M}$ generates detailed textual self-critique for each round $r$:
@@
  • Important-Word Feedback (IWF). This novel feedback approach identifies words in the input that are most important for the answer:
@@
    • Prompt-based: Following Kroeger et al. (2023) who find that LLMs can be used with high accuracy as post-hoc explainers, we prompt the model itself to assign importance scores to input words (IWF-Pmt):
@@
    • Attribution-based: Uses input feature attribution methods to quantify and assign importance scores to words with high accuracy. The implementation details are presented in Section 3.3.
@@
3.3 Attribution-Based Important-Word Feedback
While prompt-based IWF directly prompts the model to assign importance scores to input words, attribution-based IWF objectively quantifies how each input word influences the generated answer, providing interpretable and often faithful importance signals. Our approach for computing attribution-based Important-Word Feedback is illustrated in Figure 3 and provided as Algorithm 2, consisting of the following steps:
@@ 3.3 Attribution-Based Important-Word Feedback
Token-level Attribution Computation
We quantify how each token in the full model input contributes to generating each token in model output:
@@
where $a_{i,j}$ represents the attribution score of token $x_{i}$ from the full model input (prompt + task input) for the prediction of output token $y_{j}$ given all preceding context. We apply the absolute value function for two key reasons: (1) to focus on the magnitude of influence rather than its direction, as both strong positive and negative influences indicate important tokens; and (2) to prevent positive and negative attributions from canceling each other out during aggregation steps.
@@
Figure 3: Attribution-based IWF computation scheme.
@@ 4.3 Evaluation
To evaluate the faithfulness of the model generated NLEs, we employ the counterfactual test proposed by Atanasova et al. (2023).
Counterfactual Test.
The counterfactual test works by making an intervention to the original instance to get an intervened instance. If the model’s prediction changes on the intervened instance, then this intervention has influenced the model’s decision. If the NLE for the intervened instance does not mention this intervention, we consider it unfaithful, as it fails to reflect the factor that actually changed the model’s prediction.
@@
In our implementation, we adopt the random approach from Atanasova et al. (2023). Specifically, we randomly select a noun or a verb from any position in the input. For nouns, we prepend a random adjective, and for verbs, we prepend a random adverb. Our intervention generation process is entirely based on prompting GPT OpenAI et al. (2024), as this approach ensures the generation of effective, coherent, and meaningful interventions without duplications. We generate 20 unique interventions for each original instance for each dataset. The detailed intervention generation prompt and quality checks are described in Appendix C.
Unfaithfulness.
Following Atanasova et al. (2023), we employ unfaithfulness as our evaluation metric. The evaluation process consists of two steps: (1) Identify counter instances: intervened instances whose prediction changes compared to the original instance. (2) Identify unfaithful instances: counter instances whose NLEs (generated by baseline methods or SR-NLE) do not contain the intervened word (determined by string matching). The unfaithfulness rate is calculated as:
This metric allows us to directly compare the faithfulness of NLEs generated by different methods, with lower rates indicating higher faithfulness. We apply this metric consistently across all baseline methods and SR-NLE variants to ensure fair comparison.
@@ 4.4 Baselines
We compare our SR-NLE framework against two internal baselines as well as an external prior method:
@@
The initial NLEs generated by the model without any refinement process. This corresponds to $e^{0}$ in our framework and represents the typical approach used in most NLEs generation scenarios.
@@
NLEs generated using the Self-Consistency method Wang et al. (2023b), where we sample multiple explanations with temperature sampling and select the most representative explanation using the semantic centroid voting (Algorithm 3). This approach encodes all candidates using SentenceBERT Reimers and Gurevych (2019), computes their centroid in the embedding space, and selects the explanation with highest cosine similarity to this centroid. This effectively identifies the explanation that best represents the consensus meaning across all samples. This baseline represents a strong ensemble-based alternative that does not require iterative refinement. The specific configuration of sampling parameters is discussed in Section 4.5.
@@
In addition to these internal baselines, we also compare our SR-NLE with G-TEX (Yuan et al., 2025), a recent state-of-the-art method that also aims to improve explanation faithfulness. While we do not implement their approach, we report their results from the original paper for reference.

### v2

@@ Abstract
With the rapid development of Large Language Models (LLMs), Natural Language Explanations (NLEs) have become increasingly important for understanding model predictions. However, these explanations often fail to faithfully represent the model’s actual reasoning process. While existing work has demonstrated that LLMs can self-critique and refine their initial outputs for various tasks, this capability remains unexplored for improving explanation faithfulness. To address this gap, we introduce Self-critique and Refinement for Natural Language Explanations (SR-NLE), a framework that enables models to improve the faithfulness of their own explanations – specifically, post-hoc NLEs – through an iterative critique and refinement process without external supervision. Our framework leverages different feedback mechanisms to guide the refinement process, including natural language self-feedback and, notably, a novel feedback approach based on feature attribution that highlights important input words. Our experiments across three datasets and four state-of-the-art LLMs demonstrate that SR-NLE significantly reduces unfaithfulness rates, with our best method achieving an average unfaithfulness rate of 36.02%, compared to 54.81% for baseline – an absolute reduction of 18.79%. These findings reveal that the investigated LLMs can indeed refine their explanations to better reflect their actual reasoning process, requiring only appropriate guidance through feedback without additional training or fine-tuning. Our code is available at https://github.com/ymwangv/SR-NLE.
@@ 1 Introduction
With the rapid development of Large Language Models (LLMs), both closed-source models (OpenAI et al., 2024; Gemini et al., 2025) and open-source alternatives (Qwen et al., 2025; Grattafiori et al., 2024) have demonstrated remarkable capabilities across a wide range of Natural Language Processing (NLP) tasks. Yet, despite these advancements, understanding the reasoning behind their predictions remains a critical challenge – especially in applications demanding trust and accountability.
Natural Language Explanations (NLEs) have emerged as a promising solution by offering human-readable justifications for model predictions without requiring access to internal model mechanisms. However, ensuring their faithfulness remains a significant challenge. Recent studies have shown that NLEs generated by LLMs often fail to reflect the actual reasoning process of the model (Atanasova et al., 2023; Turpin et al., 2023; Lanham et al., 2023).
While prior work has primarily relied on changes of the model architecture or additional fine-tuning Yuan et al. (2025); Wang et al. (2023a); Atanasova et al. (2022), we instead explore whether models possess the capability to independently assess and refine their own explanations. Supporting this direction, recent studies have shown that LLMs are capable of improving their outputs through iterative self-refinement (Madaan et al., 2023; Shinn et al., 2023). Following this, we ask whether LLMs know if and when their NLEs are faithful to their own internal reasoning by providing self-critique and refining their NLEs for improved faithfulness.
Building on this idea, we propose Self-critique and Refinement for Natural Language Explanations (SR-NLE), a framework that enables models to improve the faithfulness of their own explanations through an iterative critique and refinement process without external supervision, as only the model itself has access to its internal reasoning and is therefore best positioned to assess explanation faithfulness. Our framework specifically targets post-hoc NLEs, where the explanation is generated after the model makes a prediction. Starting from an initial explanation, the model receives feedback identifying potential issues and generates a refined explanation accordingly. This process can be repeated multiple times, enabling incremental improvements.
@@
  • We introduce SR-NLE, a novel framework that enables models to improve the faithfulness of their explanations through iterative self-critique and refinement guided by different feedback mechanisms, without external assistance, architectural modifications or specialized training.
  • We propose and evaluate multiple feedback strategies for faithfulness, including natural language feedback (NLF) and a novel feedback mechanism – important word feedback (IWF), that leverages feature attribution to identify important input words in generated NLEs.
@@ 2 Related Works
Natural Language Explanations (NLEs) provide human-readable justifications for model predictions, traditionally obtained via supervised training on annotated datasets (Camburu et al., 2018; Rajani et al., 2019; Atanasova et al., 2020b). Recently, LLMs have enabled NLE generation via in-context learning (Brown et al., 2020). A prominent example is chain-of-thought reasoning (Wei et al., 2022; Kojima et al., 2022), where LLMs generate intermediate reasoning steps alongside the model prediction. In contrast, our work focuses on post-hoc NLEs, where explanations are generated after the model prediction is made. Despite these advances in NLE generation, numerous studies have identified a gap between generated NLEs and the model’s actual reasoning process (Atanasova et al., 2023; Turpin et al., 2023; Lanham et al., 2023). To quantify this faithfulness gap, researchers have proposed various automatic evaluation metrics, such as counterfactual tests (Atanasova et al., 2023; Siegel et al., 2024) and association-based measures (Wiegreffe et al., 2022; Parcalabescu and Frank, 2024). In this work, we adopt counterfactual tests (Atanasova et al., 2023) as our evaluation method, as they offer instance-level, automatic assessments of explanation faithfulness without requiring human annotations.
@@
Recent work has shown that LLMs can improve their own outputs via iterative self-refinement. The Self-Refine approach (Madaan et al., 2023) demonstrates that models can critique and revise their own outputs, leading to improved performance across a variety of tasks. Similarly, Shinn et al. (2023) explore self-reflection mechanisms for agent-level reasoning. While these works establish the general potential of self-improvement, they do not specifically address the challenge of improving explanation faithfulness. Building on the general idea of self-refinement, Cross-Refine (Wang et al., 2024) applies this paradigm to NLE generation. Their framework adopts a cross-model design, where one LLM generates the initial explanation and another, separate LLM provides feedback and suggestions for revision. In contrast, SR-NLE operates entirely within a single model, leveraging its internal capabilities for both critique and refinement. Furthermore, SR-NLE focuses specifically on improving faithfulness using automated counterfactual tests for objective evaluation. In comparison, Cross-Refine primarily evaluates explanation quality through multiple automated metrics, while relying on human judgments for faithfulness assessment.
@@
Input feature attribution methods quantify how much each input feature contributes to a model’s prediction. Common approaches include Shapley values (Lundberg and Lee, 2017), integrated gradients (Sundararajan et al., 2017), and attention weights (Jain and Wallace, 2019). A newer paradigm leverages prompt-based approaches (Kroeger et al., 2023), where LLMs are prompted to directly identify influential input features. The most prevalent application of these methods is to provide post-hoc explanations, helping humans understand which parts of the input most strongly influence the model’s decision-making. Beyond interpretability, these methods have also been used to construct rationales for in-context exemplars in few-shot learning to improve task accuracy. AMPLIFY(Krishna et al., 2023) trains a proxy model and applies attribution methods to extract important words, which are then converted into rationales for few-shot exemplars. Self-AMPLIFY(Bhan et al., 2024) extends this idea by removing the proxy and computing attributions directly from LMs to obtain important words, which are likewise used as rationales for exemplar construction. In the context of NLEs, G-Tex (Yuan et al., 2025) leverages attribution-derived highlights to guide the generation of NLEs through graph encoding. In this work, we propose a novel use of attribution methods to obtain feedback for improving the faithfulness of LLM-generated NLEs. Similar to SELF-AMPLIFY (Bhan et al., 2024), we apply attribution methods directly to LLMs to extract important words and use them to construct feedback to guide the model in generating more faithful NLEs.
@@ 3 Method
In this section, we present SR-NLE, a framework for improving the faithfulness of NLEs generated by LLMs. SR-NLE employs an iterative self-critique and refinement process, enabling LLMs to progressively identify faithfulness issues in their own NLEs and make targeted improvements thereof. This framework leverages the in-context learning and self-improvement capabilities of LLMs, without requiring human involvement or additional models for feedback.
@@ 3.1 Preliminary
The SR-NLE framework operates on the assumption that LLMs have the capability to identify and improve their own explanations when guided with appropriate prompts. Our framework relies entirely on a single model $\mathcal{M}$ for all components, without requiring human involvement or additional models. For an input $x$, the model first predicts an answer $y$ and produces an initial explanation $e^{0}$, then, through an iterative process of self-critique and refinement, after each round $r$, we obtain a progressively improved explanation $e^{r}$. To direct the model in different stages of the framework, we employ four categories of prompts: $p_{\text{ans}}$ for answer generation, $p_{\text{exp}}$ for explanation generation, $p_{\text{fb}}$ for feedback generation, and $p_{\text{ref}}$ for refinement generation, where both feedback and refinement prompts have two variants corresponding to our two feedback approaches: natural language feedback (NLF) and important word feedback (IWF). Throughout the framework, we use $\oplus$” to denote filling a prompt template with its variables.
Figure 2: The SR-NLE framework. (a) Answer and Explanation Generation Phase: The framework produces the answer $x$ and initial explanation $e^{0}$. (b) Iterative Critique and Refinement Phase: The framework iteratively improves explanations through feedback-refinement loops over multiple rounds.
@@ 3.2 SR-NLE Framework
Our SR-NLE framework consists of two main phases: (a) Answer and Explanation Generation, which produces the answer and its initial explanation, and (b) Iterative Critique and Refinement, which progressively improves the explanation through multiple rounds. Figure 2 illustrates this two-phase process. The algorithmic formulation is provided as Algorithm 1.
@@ 3.2.2 Iterative Critique and Refinement
For each refinement round $r$, the model generates feedback on the preceding explanation $e^{r-1}$. We explore two distinct feedback approaches:
  • Natural Language Feedback (NLF). With this approach, $\mathcal{M}$ generates detailed textual self-critique for each round $r$:
@@
  • Important Word Feedback (IWF). This novel feedback approach leverages attribution explanations, which mark specific input tokens DeYoung et al. (2020) or segments Ray Choudhury et al. (2023) critical to a model’s prediction. While these explanations may lack the plausibility of NLEs (Jie et al., 2024), their faithfulness is straightforward to measure and has seen significant improvements (Sun et al., 2025; Atanasova et al., 2020a). We hypothesise that such explanations can enhance NLE faithfulness by providing explicit feedback about which input elements should be emphasized in the generated explanation. Our approach identifies words in the input that are most important for the answer:
@@
    • Prompt-based: Following Kroeger et al. (2023), who find that LLMs can be used with high accuracy as post-hoc explainers, we prompt the model itself to assign importance scores to input words (IWF-Pmt):
@@
    • Attribution-based: Following Bhan et al. (2024), we use input feature attribution methods to quantify and assign importance scores to words (IWF-Attr). We detail our method for computing the IWF-Attr Score in Section $3.$
@@
3.3 Attribution-Based IWF Score
While prompt-based IWF directly prompts the model to assign importance scores to input words, attribution-based IWF computes these scores using feature attribution methods. Our approach for computing the attribution-based IWF Score is illustrated in Figure 3 and detailed in Algorithm 2, consisting of the following steps:
@@ 3.3 Attribution-Based IWF Score
Token-level Computation
We quantify how each token in the full model input contributes to generating each token in the model output:
@@
where $a_{i,j}$ represents the attribution score of token $x_{i}$ from the full model input (prompt + task input) for the prediction of output token $y_{j}$ given all preceding context. We apply the absolute value function for two key reasons: (1) to focus on the magnitude of influence rather than its direction, as both strong positive and negative influences indicate important tokens; and (2) to prevent positive and negative attributions from cancelling each other out during aggregation steps.
@@
Figure 3: Illustration of attribution-based IWF Score.
@@ 4.3 Evaluation
To evaluate the faithfulness of the model-generated NLEs, we employ the counterfactual test proposed by Atanasova et al. (2023). The counterfactual test works by making an intervention to the original instance to get an intervened instance. The evaluation then consists of two steps: (1) Identify counter instances: intervened instances whose prediction changes compared to the original instance. (2) Identify unfaithful instances: counter instances whose NLEs (generated by baseline methods or SR-NLE) do not contain the intervened word (determined by string matching). The unfaithfulness rate is calculated as:
This metric allows us to directly compare the faithfulness of NLEs generated by different methods, with lower rates of unfaithfulness being more desirable. We apply this metric consistently across all baseline methods and SR-NLE variants to ensure fair comparison.
@@
In our implementation, we adopt the random approach from Atanasova et al. (2023). Specifically, we randomly select a noun or a verb from any position in the input. For nouns, we prepend a random adjective, and for verbs, we prepend a random adverb. Different from Atanasova et al. (2023), we further employ prompting GPT-4o OpenAI et al. (2024), to ensure the generation of multiple effective, coherent, and meaningful interventions for the same instance without duplications. We generate 20 unique interventions for each original instance from each dataset. The detailed intervention generation prompt and quality checks are described in Appendix C.
@@ 4.4 Baselines
We compare our SR-NLE framework against two baselines suggested by us, as well as an existing prior method:
@@
The initial NLEs were generated by the model without any refinement process. This corresponds to $e^{0}$ in our framework and represents the typical approach used in most NLE generation scenarios.
@@
NLEs generated using the Self-Consistency method Wang et al. (2023b), where we sample multiple explanations with temperature sampling and select the most representative explanation using the semantic centroid voting (Algorithm 3). This approach encodes all candidates using SentenceBERT Reimers and Gurevych (2019), computes their centroid in the embedding space, and selects the explanation with the highest cosine similarity to this centroid. This effectively identifies the explanation that best represents the consensus meaning across all samples. This baseline represents a strong ensemble-based alternative that does not require iterative refinement. The specific configuration of sampling parameters is discussed in Section 4.5.
@@
We also compare our SR-NLE with G-TEX (Yuan et al., 2025), a recent state-of-the-art method that also aims to improve explanation faithfulness. While we do not implement their approach, we report their results from the original paper for reference.